# Functional normalizing flow for statistical inverse problems of partial differential equations

### v1

@@ 2.3.2 Functional Projected Transformation Flow
The second model is termed functional projected transformation flow. Let $\{\phi_{i}\}_{i=1}^{M}$ be the first $M$ eigenvectors of the prior covariance operator, and we define two operators $\mathcal{P}$, $\mathcal{Q}$:
@@
Let $\mathcal{H}_{u}$ be a separable Hilbert space equipped with the Gaussian measure $\mu_{0}=\mathcal{N}(0,\mathcal{C})$, and $\mathcal{H}=\mathcal{H}(\mu_{0})$ be the Cameron-Martin space of the measure $\mu_{0}$. For each $n=1,2,\ldots,N$, let $f_{\theta_{n}}^{(n)}(u)=u+\mathcal{Q}R_{n}(\mathcal{P}u+b_{n})$, where $\mathcal{F}_{\theta_{n}}^{(n)}$ is an operator from $\mathcal{H}_{u}$ to $\mathcal{H}_{u}$ and $\theta_{n}=\{R_{n},b_{n}\}$ be its parameters. The composite transformation $f_{\theta}$ is defined by $f_{\theta}(u)=f_{\theta_{N}}^{(N)}\circ f_{\theta_{N-1}}^{(N-1)}\circ\cdots\circ f_{\theta_{1}}^{(1)}(u)$, and $\mu_{f_{\theta}}$ denotes its corresponding push-forward measure given by $\mu_{f_{\theta}}=\mu_{0}\circ f_{\theta}^{-1}$. Assuming the following conditions hold for all $n=1,2,\ldots,N$ and any value of $\theta_{n}$:
@@
Then we have $\mu_{f_{\theta}}\sim\mu$ and
where $u_{n}=f_{\theta_{n}}^{(n)}\circ f_{\theta_{n-1}}^{(n-1)}\circ\cdots\circ f_{\theta_{1}}^{(1)}(u)$ for each $n=1,2,\ldots,N$.
The point spectrum of $I+\mathcal{Q}R_{n}\mathcal{P}$ corresponds one-to-one with the eigenvalues of $I_{M}+R_{n}$. To ensure that the eigenvalues of $R_{n}$ are strictly greater than $-1$, we restrict $R_{n}$ to be an upper triangular matrix with diagonal elements in $(-1,+\infty)$.
@@ 2.4 Nonlinear Transformation
Let $\mathcal{H}_{u}$ be a separable Hilbert space, $f_{\theta_{n}}^{(n)}:\mathcal{H}_{u}\rightarrow\mathcal{H}_{u}$, and $f_{\theta_{n}}^{(n)}(u)=u+\mathcal{F}_{\theta_{n}}^{(n)}(u)$. Assuming the following conditions hold for all $n=1,2,\ldots,N$ and any value of $\theta_{n}$:
@@
    The operator $\mathcal{F}_{\theta_{n}}^{(n)}$ is a finite rank operator.
@@
    The image space of operator $\mathcal{F}_{\theta_{n}}^{(n)}$ is bounded in $\mathcal{H}_{u}$.
@@ 2.4.1 Functional Planar Flow
The first nonlinear model, termed functional planar flow, is a generalization of the planar flow in Euclidean space [40]. Each layer of functional Householder flow employs the transformation:
@@
where $u_{n},w_{n}\in\mathcal{H}_{u}$, $b_{n}\in\mathbb{R}$, $h(x)=tanh(x)$ and the parameters set is given by $\theta_{n}=\{u_{n},w_{n},b_{n}\}$. The following theorem guarantees the model’s adherence to the theoretical framework proposed in Subsection 2.2.
@@
Let $\mathcal{H}_{u}$ be a separable Hilbert space equipped with the Gaussian measure $\mu_{0}=\mathcal{N}(0,\mathcal{C})$, and $\mathcal{H}=\mathcal{H}(\mu_{0})$ be the Cameron-Martin space of the measure $\mu_{0}$. For each $n=1,2,\ldots,N$, let $f_{\theta_{n}}^{(n)}(u)=u+u_{n}h(\langle w_{n},u\rangle_{\mathcal{H}_{u}}+b_{n})$, where $\mathcal{F}_{\theta_{n}}^{(n)}$ is an operator from $\mathcal{H}_{u}$ to $\mathcal{H}_{u}$ and $\theta_{n}=\{v_{n},b_{n}\}$ be its parameters. The composite transformation $f_{\theta}$ is defined by $f_{\theta}(u)=f_{\theta_{N}}^{(N)}\circ f_{\theta_{N-1}}^{(N-1)}\circ\cdots\circ f_{\theta_{1}}^{(1)}(u)$, and $\mu_{f_{\theta}}$ denotes its corresponding push-forward measure given by $\mu_{f_{\theta}}=\mu_{0}\circ f_{\theta}^{-1}$. Assuming the following conditions hold for all $n=1,2,\ldots,N$ and any value of $\theta_{n}$:
@@
Then we have $\mu_{f_{\theta}}\sim\mu$ and
where $u_{n}=f_{\theta_{n}}^{(n)}\circ f_{\theta_{n-1}}^{(n-1)}\circ\cdots\circ f_{\theta_{1}}^{(1)}(u)$ for each $n=1,2,\ldots,N$.
Given that the parameters $u_{n}$ and $w_{n}$ of our flow model are functions, efficient parameterization is paramount to both computational efficiency and ensuring compliance with the theorem’s constraints. To address this, we propose a novel parameterization method for the flow model. Let $\{\phi_{i}\}_{i=1}^{M}$ represent the first $M$ eigenvectors of the prior covariance operator. Our parameterization of $u_{n}$ and $w_{n}$ involves a two-step process. Firstly, auxiliary vectors $\hat{u}_{n}$, $\hat{w}_{n}$ are introduced and parameterized as
@@
Like the functional Householder flow, the image of $\mathcal{F}_{\theta_{n}}^{(n)}$ is constrained to a one-dimensional space, which limits the model’s expressive capabilities. To effectively approximate target measures that are significantly different from the prior, a deep network architecture with multiple layers is often necessary. The effectiveness of this approach is demonstrated through numerical experiments in Section 4.
@@ 2.4.2 Functional Sylvester Flow
The second nonlinear model we are exploring, known as functional Sylvester flow, is an extension of Sylvester flow, a concept traditionally applied in Euclidean space [46]. Each layer of functional Sylvester flow utilizes the following transformation:
@@
where $\mathcal{A}_{n}$ is a bounded linear operator mapping from $\mathcal{H}_{u}$ to $\mathbb{R}^{M}$, $\mathcal{B}_{n}$ is a bounded linear operator mapping from $\mathbb{R}^{M}$ to $\mathcal{H}_{u}$, $b_{n}\in\mathbb{R}^{M}$, $h(x)=tanh(x)$, and the parameter set is given by $\theta_{n}=\{\mathcal{A}_{n},\mathcal{B}_{n},b_{n}\}$. The following theorem ensures that this model adheres to the theoretical framework we outlined in Subsection 2.2.
@@
Let $\mathcal{H}_{u}$ be a separable Hilbert space equipped with the Gaussian measure $\mu_{0}=\mathcal{N}(0,\mathcal{C})$, and $\mathcal{H}=\mathcal{H}(\mu_{0})$ be the Cameron-Martin space of the measure $\mu_{0}$. For each $n=1,2,\ldots,N$, let $f_{\theta_{n}}^{(n)}(u)=u+\mathcal{A}_{n}h(\mathcal{B}_{n}u+b_{n})$, where $\mathcal{F}_{\theta_{n}}^{(n)}$ is an operator from $\mathcal{H}_{u}$ to $\mathcal{H}_{u}$ and $\theta_{n}=\{\mathcal{A}_{n},\mathcal{B}_{n},b_{n}\}$ be its parameters. The composite transformation $f_{\theta}$ is defined by $f_{\theta}(u)=f_{\theta_{N}}^{(N)}\circ f_{\theta_{N-1}}^{(N-1)}\circ\cdots\circ f_{\theta_{1}}^{(1)}(u)$, and $\mu_{f_{\theta}}$ denotes its corresponding push-forward measure given by $\mu_{f_{\theta}}=\mu_{0}\circ f_{\theta}^{-1}$. Assuming the following conditions hold for all $n=1,2,\ldots,N$ and any value of $\theta_{n}$:
@@
    The operator $\mathcal{A}_{n}$ is a finite rank operator.
@@
Then we have $\mu_{f_{\theta}}\sim\mu$ and
where $u_{n}=f_{\theta_{n}}^{(n)}\circ f_{\theta_{n-1}}^{(n-1)}\circ\cdots\circ f_{\theta_{1}}^{(1)}(u)$ for any $n=1,2,\ldots,N$.
Given that the parameters $\mathcal{A}_{n}$ and $\mathcal{B}_{n}$ of the flow model are operators, efficient parameterization is essential for both computational efficiency and ensuring compliance with the theorem’s constraints. We select the first $M$ eigenvectors of the prior covariance operator, denoted as $\{\phi_{i}\}_{i=1}^{M}$. The parameterization of $\mathcal{B}_{n}$ involves two steps:
@@
    The first step is defined by the operator $\mathcal{B}_{n}^{(1)}$ as follows:
    $p_{\mathcal{B}}=\mathcal{B}_{n}^{(1)}u:=\begin{pmatrix}\langle u,\phi_{1}\rangle_{\mathcal{H}_{u}}\\ \langle u,\phi_{2}\rangle_{\mathcal{H}_{u}}\\ \vdots\\ \langle u,\phi_{M}\rangle_{\mathcal{H}_{u}}\\ \end{pmatrix}.$
@@
    The second step is defined by the operator $\mathcal{B}_{n}^{(2)}$ as follows:
    $\mathcal{B}_{n}^{(2)}p_{\mathcal{B}}:=R_{\mathcal{B}}^{n}p_{\mathcal{B}},$ (2.14)
@@
    We map the vector $d$ to $p_{\mathcal{A}}\in\mathbb{R}^{M}$ by a square matrix $R_{\mathcal{A}}^{n}$. This step is defined by the operator $\mathcal{A}_{n}^{(1)}$ as follows:
    $p_{\mathcal{A}}=\mathcal{A}_{n}^{(1)}d:=R_{\mathcal{A}}^{n}d.$ (2.15)
@@
    The second step is defined as follows:
    $\mathcal{A}_{n}^{(2)}p_{\mathcal{A}}:=\sum\limits_{i=1}^{M}p_{\mathcal{A}}^{i}\phi_{i},$
@@
Let $\mathcal{H}_{u}$ be a separable Hilbert space, $\mathcal{B}$ be a bounded linear operator from $\mathcal{H}_{u}$ to $\mathbb{R}^{M}$, $\mathcal{A}$ be a bounded linear operator from $\mathbb{R}^{M}$ to $\mathcal{H}_{u}$. We have:
@@
${\det}_{1}(I+\mathcal{A}_{n}\mathrm{diag}(h^{{}^{\prime}}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n})=\det(I_{M}+\mathrm{diag}(h^{{}^{\prime}}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n}\mathcal{A}_{n}).$
This allows us to efficiently compute the Fredholm determinant of the operator $Df_{\theta_{n}}^{(n)}(u)=I+\mathcal{A}_{n}\mathrm{diag}(h^{{}^{\prime}}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n}$. Theorem 2.9 establishes the equivalence between the point spectrum of $I+\mathcal{A}_{n}\mathrm{diag}(h^{{}^{\prime}}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n}$ and the eigenvalues of the matrix $I_{M}+\mathrm{diag}(h^{{}^{\prime}}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n}\mathcal{A}_{n}$. In order to ensure that the eigenvalues of $\mathrm{diag}(h^{{}^{\prime}}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n}\mathcal{A}_{n}$ are not in $(-\infty,-1]$, we set $R_{\mathcal{B}}^{n}$ in (2.14) as an upper triangular matrix, $R_{\mathcal{A}}^{n}$ in (2.15) as a lower triangular matrix, the diagonal elements of $R_{\mathcal{B}}^{n}$ are all 1, and the diagonal elements of $R_{\mathcal{A}}^{n}$ are all greater than $-1$. Through this method, $\mathrm{diag}(h^{{}^{\prime}}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n}\mathcal{A}_{n}$ is an upper triangular matrix, and the diagonal elements are all greater than $-1$. Hence we can compute the determinant of $I_{M}+\mathrm{diag}(h^{{}^{\prime}}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n}\mathcal{A}_{n}$ quickly.
@@ 2.5 Discretization Invariance
The model, defined within a function space, inherently implies discretization invariance. In general, it can be summarized as the following three properties [29]:
@@
In this subsection, we investigate the discretization invariance of the operator $\mathcal{F}_{\theta_{n}}^{(n)}$ within the flow model $f_{\theta_{n}}^{(n)}=I+\mathcal{F}_{\theta_{n}}^{(n)}$. Following the approach outlined in [29], a formal definition of the discretization invariance is provided.
@@
$D\subset\bigcup_{x\in D_{L}}\{y\in\mathbb{R}^{d}:\lVert y-x\rVert<\epsilon\}.$
@@
Suppose $\mathcal{H}_{u}$ is a Hilbert space of $\mathbb{R}^{m}$ valued functions on the domain $D\subset\mathbb{R}^{d}$. For any $\theta_{n}$, let $\mathcal{F}_{\theta_{n}}^{(n)}:\mathcal{H}_{u}\rightarrow\mathcal{H}_{u}$ be an operator, $D_{L}$ be an $L$-point discretization of $D$, and operator ^$\hat{\mathcal{F}}_{\theta_{n}(L)}^{(n)}:\mathbb{R}^{Lm}\rightarrow\mathcal{H}_{u}$. For any $K\subset C(D)$ compact, we define the discretized uniform risk as
@@
Given $a$ discrete refnement ${\{D_{n}\}}_{n=1}^{\infty}$ of the domain $D\subset\mathbb{R}^{d}$. For any fixed $\theta_{n}$, we say $\mathcal{F}_{\theta_{n}}^{(n)}$ is discretization-invariant if there exists a sequence of maps ^$\hat{\mathcal{F}}_{\theta_{n}(1)}^{(n)},\hat{\mathcal{F}}_{\theta_{n}(2)}^{(n)},\ldots$ where ^$\hat{\mathcal{F}}_{\theta_{n}(L)}^{(n)}:\mathbb{R}^{Lm}\to\mathcal{H}_{u}$ such that, for any compact set $K\subset C(D)$,
@@
Let $D\subset\mathbb{R}^{d}$ be a domain for some $d\in\mathbb{N}$. Suppose that $\mathcal{H}_{u}$ can be continuously embedded in $C(D)$. Then for any $n\in\mathbb{N}$, the layers $\mathcal{F}_{\theta_{n}}^{(n)}:\mathcal{H}_{u}\rightarrow\mathcal{H}_{u}$ of functional planar flow, functional Householder flow, functional Sylvester flow, and functional projected transformation flow are discretization-invariant.
Next, we compare and contrast the functional normalizing flow framework with standard normalizing flow in Euclidean space. A key property contributing to the success of normalizing flow in computer vision is the tractability of the density of the post-transformed distribution. To achieve this, the transformation $\widetilde{f}(z)$ of normalizing flow in Euclidean space must be invertible [40, 18]. In this case, the post-transformed density can be expressed as
@@
where $z^{\prime}=\widetilde{f}(z)$, $z,z^{\prime}\in\mathbb{R}^{L}$ and $\widetilde{f}$ maps from $\mathbb{R}^{L}$ to $\mathbb{R}^{L}$, the densities of the post-transformed and pre-transformed distributions are denoted by $p_{1}(z)$ and $p_{0}(z)$, respectively. This tractable form allows for the computation of the Kullback-Leibler divergence between the target distribution and the post-transformed distribution, which is used as the loss function.
In this paper, we demonstrate the applicability of similar ideas within the infinite-dimensional Hilbert space $\mathcal{H}_{u}$, aligning well with the Bayesianize-then-Discretize approach for inverse problem. Similar to the normalizing flow in Euclidean space, ensuring the invertibility of the transformation $f_{\theta}$ is crucial for the tractability of the RN derivative between the post-transformed and pre-transformed measures [7]. While various different approaches have been proposed to ensure the invertibility of normalizing flow transformations in Euclidean space [40, 18], we present general theorems (Lemmas 2.1 and 2.2) that establish conditions for the invertibility of transformations within the infinite-dimensional function space. This ensures the tractability of the RN derivative between the post-transformed measure and the prior, enabling the calculation of the KL divergence as the loss function.
@@
where $\bm{u}_{n}=(u_{n}^{(1)},\ldots,u_{n}^{(L)})$, $\bm{w}_{n}=(\frac{w_{n}^{(1)}}{L},\ldots,\frac{w_{n}^{(L)}}{L})$, $b_{n}\in\mathbb{R}$ is constant, $h(x)=tanh(x)$. Parameters $\theta_{n}$ can be regard as $\theta_{n}=\{\bm{u}_{n},\bm{w}_{n},b_{n}\}$. The construction (2.16) can be generalized to arbitrary number of layers by composing several transformations together
$\widetilde{f}_{\theta}(\widetilde{u})=\widetilde{f}_{\theta_{N}}^{(N)}\circ\widetilde{f}_{\theta_{N-1}}^{(N-1)}\circ\cdots\circ\widetilde{f}_{\theta_{1}}^{(1)}(\widetilde{u}).$ (2.18)
The obtained model (2.18) is the planar flow defined in the Euclidean space [40]. It is evident that when employing the planar flow (2.17) in function spaces, the parameterization of $w_{n}$ becomes contingent upon the specific discretization of the input $\bm{u}$. Likewise, the parameterization of $u_{n}$ depends on the desired discretization of the output. Consequently, for varying discretization requirements, the entire model necessitates retraining, precluding the ability to generalize to arbitrary discretized inputs and outputs. Consequently, standard planar flow in Euclidean space are not consistent within the infinite-dimensional function space, and it is not discretization-invariant.

### v2

@@ 2.4.2 Functional Projected Transformation Flow
The second model is termed functional projected transformation flow. Let $\{\phi_{i}\}_{i=1}^{M}$ be the first $M$ eigenfunctions of the prior covariance operator, and we define two operators $\mathcal{P}$, $\mathcal{Q}$:
@@
Let $f_{\theta_{n}}^{(n)}(u)=u+\mathcal{Q}R_{n}(\mathcal{P}u+b_{n})$. Assume that the following conditions hold for all $n=1,2,\ldots,N$:
@@
Then $f_{\theta_{n}}^{(n)}$ satisfies all the conditions of Theorem 2.3.
@@ 2.5 Nonlinear Transformation
Let $\mathcal{H}_{u}$ be a separable Hilbert space, $f_{\theta_{n}}^{(n)}\mathrel{\mathop{\ordinarycolon}}\mathcal{H}_{u}\rightarrow\mathcal{H}_{u}$, and $f_{\theta_{n}}^{(n)}(u)=u+\mathcal{F}_{\theta_{n}}^{(n)}(u)$. Assume that the following conditions hold for all $n=1,2,\ldots,N$:
@@
    The operator $\mathcal{F}_{\theta_{n}}^{(n)}$ has finite dimensional range.
@@
    $\mathcal{F}_{\theta_{n}}^{(n)}(\mathcal{H}_{u})$ is a bounded subset in $\mathcal{H}_{u}$.
@@ 2.5.1 Functional Planar Flow
The first nonlinear model, termed functional planar flow, is a generalization of the planar flow in Euclidean space planarflow. Each layer of functional planar flow employs the transformation:
@@
where $u_{n},w_{n}\in\mathcal{H}_{u}$, $b_{n}\in\mathbb{R}$, $h(x)=\tanh(x)$ and the parameters set is given by $\theta_{n}=\{u_{n},w_{n},b_{n}\}$. The following theorem guarantees the model’s adherence to the theoretical framework proposed in Subsection 2.2.
@@
Let $f_{\theta_{n}}^{(n)}(u)=u+u_{n}h(\langle w_{n},u\rangle_{\mathcal{H}_{u}}+b_{n})$. Assume that the following conditions hold for all $n=1,2,\ldots,N$:
@@
Then $f_{\theta_{n}}^{(n)}$ satisfies all the conditions of Theorem 2.3.
Given that the parameters $u_{n}$ and $w_{n}$ of our flow model are functions, an efficient parameterization is essential both for computational efficiency and for ensuring compliance with the assumptions of the theorem. To address this, we propose a novel parameterization method for the flow model. Let $\{\phi_{i}\}_{i=1}^{M}$ represent the first $M$ eigenfunctions of the prior covariance operator. Our parameterization of $u_{n}$ and $w_{n}$ involves a two-step process. Firstly, auxiliary functions $\hat{u}_{n}$, $\hat{w}_{n}$ are introduced and parameterized as
@@
Similar with functional Householder flow, the image of $\mathcal{F}_{\theta_{n}}^{(n)}$ is constrained to a one-dimensional space, which limits the model’s expressive capabilities. To effectively approximate target measures that are significantly different from the prior, a deep network architecture with multiple layers is often necessary. The effectiveness of this approach is demonstrated through numerical experiments in Section 4.
@@ 2.5.2 Functional Sylvester Flow
The second nonlinear model we are exploring, known as functional Sylvester flow, is an extension of Sylvester flow, a concept traditionally applied in Euclidean space sylvesterflow. Each layer of functional Sylvester flow utilizes the following transformation:
@@
where $\mathcal{B}_{n}$ is a bounded linear operator mapping from $\mathcal{H}_{u}$ to $\mathbb{R}^{M}$, $\mathcal{A}_{n}$ is a bounded linear operator mapping from $\mathbb{R}^{M}$ to $\mathcal{H}_{u}$, $b_{n}\in\mathbb{R}^{M}$, $h(x)=\tanh(x)$, and the parameters set is given by $\theta_{n}=\{\mathcal{A}_{n},\mathcal{B}_{n},b_{n}\}$. The following theorem ensures that this model adheres to the theoretical framework we outlined in Subsection 2.2.
@@
Let $f_{\theta_{n}}^{(n)}(u)=u+\mathcal{A}_{n}h(\mathcal{B}_{n}u+b_{n})$. Assume that the following conditions hold for all $n=1,2,\ldots,N$:
@@
    The operator $\mathcal{A}_{n}$ has finite dimensional range.
@@
Then $f_{\theta_{n}}^{(n)}$ satisfies all the conditions of Theorem 2.3.
Given that the parameters $\mathcal{A}_{n}$ and $\mathcal{B}_{n}$ of the flow model are operators, efficient parameterization is essential for both computational efficiency and ensuring compliance with the theorem’s constraints. We select the first $M$ eigenfunctions of the prior covariance operator, denoted as $\{\phi_{i}\}_{i=1}^{M}$. The parameterization of $\mathcal{B}_{n}$ involves two steps:
@@
    First, $\mathcal{B}_{n}^{(1)}$ projects $u$ onto a finite-dimensional subspace, defining $p_{\mathcal{B}}$:
    $p_{\mathcal{B}}=\mathcal{B}_{n}^{(1)}u\mathrel{\mathop{\ordinarycolon}}=\begin{pmatrix}\langle u,\phi_{1}\rangle_{\mathcal{H}_{u}}\\ \langle u,\phi_{2}\rangle_{\mathcal{H}_{u}}\\ \vdots\\ \langle u,\phi_{M}\rangle_{\mathcal{H}_{u}}\end{pmatrix}$.
@@
    Second, $\mathcal{B}_{n}^{(2)}$ applies a linear transformation on $p_{\mathcal{B}}$ via the square matrix $R_{\mathcal{B}}^{n}$:
    $\mathcal{B}_{n}^{(2)}p_{\mathcal{B}}\mathrel{\mathop{\ordinarycolon}}=R_{\mathcal{B}}^{n}p_{\mathcal{B}},$ (2.14)
@@
    The first step, $\mathcal{A}_{n}^{(1)}$ maps the vector $d$ to $p_{\mathcal{A}}\in\mathbb{R}^{M}$ using the square matrix $R_{\mathcal{A}}^{n}$:
    $p_{\mathcal{A}}=\mathcal{A}_{n}^{(1)}d\mathrel{\mathop{\ordinarycolon}}=R_{\mathcal{A}}^{n}d.$ (2.15)
@@
    The second step, $\mathcal{A}_{n}^{(2)}$ maps the vector $p_{\mathcal{A}}$ back to the function space spanned by the basis functions $\{\phi_{i}\}$:
    $\mathcal{A}_{n}^{(2)}p_{\mathcal{A}}\mathrel{\mathop{\ordinarycolon}}=\sum\limits_{i=1}^{M}p_{\mathcal{A}}^{i}\phi_{i},$
@@
Let $\mathcal{H}_{u}$ be a separable Hilbert space. Let $\mathcal{B}\mathrel{\mathop{\ordinarycolon}}\mathcal{H}_{u}\to\mathbb{R}^{M}$ and $\mathcal{A}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{M}\to\mathcal{H}_{u}$ be bounded linear operators. Then the following statements hold:
@@
${\det}_{1}(I+\mathcal{A}_{n}\mathrm{diag}(h^{\prime}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n})=\det(I_{M}+\mathrm{diag}(h^{\prime}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n}\mathcal{A}_{n}).$
This allows us to efficiently compute the Fredholm determinant of the operator $Df_{\theta_{n}}^{(n)}(u)=I+\mathcal{A}_{n}\mathrm{diag}(h^{\prime}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n}$. Theorem 2.8 establishes the equivalence between the point spectrum of $I+\mathcal{A}_{n}\mathrm{diag}(h^{\prime}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n}$ and the eigenvalues of the matrix $I_{M}+\mathrm{diag}(h^{\prime}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n}\mathcal{A}_{n}$. In order to ensure that the eigenvalues of $\mathrm{diag}(h^{\prime}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n}\mathcal{A}_{n}$ are not in $(-\infty,-1]$, we set $R_{\mathcal{B}}^{n}$ in (2.14) as an upper triangular matrix, $R_{\mathcal{A}}^{n}$ in (2.15) as a lower triangular matrix, the diagonal elements of $R_{\mathcal{B}}^{n}$ are all 1, and the diagonal elements of $R_{\mathcal{A}}^{n}$ are all greater than 1-1. Through this method, $\mathrm{diag}(h^{\prime}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n}\mathcal{A}_{n}$ is an upper triangular matrix, and the diagonal elements are all greater than 1-1. Hence we can compute the determinant of $I_{M}+\mathrm{diag}(h^{\prime}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n}\mathcal{A}_{n}$ quickly.
@@ 2.6 Discretization Invariance
The model, defined within a function space, inherently implies discretization invariance. In general, it can be summarized as the following three properties kovachki2023neural:
@@
In this subsection, we investigate the discretization invariance of the operator $\mathcal{F}_{\theta_{n}}^{(n)}$ within the flow model $f_{\theta_{n}}^{(n)}=I+\mathcal{F}_{\theta_{n}}^{(n)}$. Following the approach outlined in kovachki2023neural, a formal definition of the discretization invariance is provided.
@@
$D\subset\bigcup_{x\in D_{L}}\{y\in\mathbb{R}^{d}\mathrel{\mathop{\ordinarycolon}}\lVert y-x\rVert<\epsilon\}.$
@@
Suppose $\mathcal{H}_{u}$ is a Hilbert space of $\mathbb{R}^{m}$ valued functions on the domain $D\subset\mathbb{R}^{d}$. For any $\theta_{n}$, let $\mathcal{F}_{\theta_{n}}^{(n)}\mathrel{\mathop{\ordinarycolon}}\mathcal{H}_{u}\rightarrow\mathcal{H}_{u}$ be an operator, $D_{L}$ be an $L$-point discretization of $D$, and operator $\hat{\mathcal{F}}_{\theta_{n}(L)}^{(n)}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{Lm}\rightarrow\mathcal{H}_{u}$. For any $K\subset C(D)$ compact, we define the discretized uniform risk as
@@
Given $a$ discrete refnement ${\{D_{n}\}}_{n=1}^{\infty}$ of the domain $D\subset\mathbb{R}^{d}$. For any fixed $\theta_{n}$, we say $\mathcal{F}_{\theta_{n}}^{(n)}$ is discretization-invariant if there exists a sequence of maps $\hat{\mathcal{F}}_{\theta_{n}(1)}^{(n)},\hat{\mathcal{F}}_{\theta_{n}(2)}^{(n)},\ldots$ where $\hat{\mathcal{F}}_{\theta_{n}(L)}^{(n)}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{Lm}\to\mathcal{H}_{u}$ such that, for any compact set $K\subset C(D)$,
@@
Let $D\subset\mathbb{R}^{d}$ be a domain for some $d\in\mathbb{N}$. Suppose that $\mathcal{H}_{u}$ can be continuously embedded in $C(D)$. Then for any $n\in\mathbb{N}$, the layers $\mathcal{F}_{\theta_{n}}^{(n)}\mathrel{\mathop{\ordinarycolon}}\mathcal{H}_{u}\rightarrow\mathcal{H}_{u}$ of functional planar flow, functional Householder flow, functional Sylvester flow, and functional projected transformation flow are discretization-invariant.
Next, we compare and contrast the functional normalizing flow framework with standard normalizing flow in Euclidean space. A key property contributing to the success of normalizing flow in computer vision is the tractability of the density of the post-transformed distribution. To achieve this, the transformation $\widetilde{f}(z)$ of normalizing flow in Euclidean space must be invertible planarflow, nvp. In this case, the post-transformed density can be expressed as
@@
where $^{\prime}=\widetilde{f}(z), z,z$$z^{\prime}\in\mathbb{R}^{L}$ and $\widetilde{f}$ maps from $\mathbb{R}^{L}$ to $\mathbb{R}^{L}$, the densities of the post-transformed and pre-transformed distributions are denoted by $p_{1}(z)$ and $p_{0}(z)$, respectively. This tractable form allows for the computation of the Kullback-Leibler divergence between the target distribution and the post-transformed distribution, which is used as the loss function.
In this paper, we demonstrate the applicability of similar ideas within the infinite-dimensional Hilbert space $\mathcal{H}_{u}$, aligning well with the Bayesianize-then-Discretize approach for inverse problem. Similar to the normalizing flow in Euclidean space, ensuring the invertibility of the transformation $f_{\theta}$ is crucial for the tractability of the RN derivative between the post-transformed and pre-transformed measures gaussianmeasure. While various different approaches have been proposed to ensure the invertibility of normalizing flow transformations in Euclidean space planarflow, nvp, we present general theorems (Lemmas 2.1 and 2.2) that establish conditions for the invertibility of transformations within the infinite-dimensional function space. This ensures the tractability of the RN derivative between the post-transformed measure and the prior, enabling the calculation of the KL divergence as the loss function.
@@
where $\bm{u}_{n}=(u_{n}^{(1)},\ldots,u_{n}^{(L)})$, $\bm{w}_{n}=(\frac{w_{n}^{(1)}}{L},\ldots,\frac{w_{n}^{(L)}}{L})$, $b_{n}\in\mathbb{R}$ is constant, $h(x)=\tanh(x)$. Parameters $\theta_{n}$ can be regard as $\theta_{n}=\{\bm{u}_{n},\bm{w}_{n},b_{n}\}$. The construction (2.16) can be generalized to arbitrary number of layers by composing several transformations together
[…]
The obtained model (2.18) is the planar flow defined in the Euclidean space planarflow. It is evident that when employing the planar flow (2.17) in function spaces, the parameterization of $w_{n}$ becomes contingent upon the specific discretization of the input $\bm{u}$. Likewise, the parameterization of $u_{n}$ depends on the desired discretization of the output. Consequently, for varying discretization requirements, the entire model necessitates retraining, precluding the ability to generalize to arbitrary discretized inputs and outputs. Consequently, standard planar flow in Euclidean space are not consistent within the infinite-dimensional function space, and it is not discretization-invariant.